# Robust Multi-Task Learning with Excess Risks

### v2

@@ Abstract
Multi-task learning (MTL) considers learning a joint model for multiple tasks by optimizing a convex combination of all task losses. To solve the optimization problem, existing methods use an adaptive weight updating scheme, where task weights are dynamically adjusted based on their respective losses to prioritize difficult tasks. However, these algorithms face a great challenge whenever label noise is present, in which case excessive weights tend to be assigned to noisy tasks that have relatively large Bayes optimal errors, thereby overshadowing other tasks and causing performance to drop across the board. To overcome this limitation, we propose Multi-Task Learning with Excess Risks (ExcessMTL), an excess risk-based task balancing method that updates the task weights by their distances to convergence instead. Intuitively, ExcessMTL assigns higher weights to worse-trained tasks that are further from convergence. To estimate the excess risks, we develop an efficient and accurate method with Taylor approximation. Theoretically, we show that our proposed algorithm achieves convergence guarantees and Pareto stationarity. Empirically, we evaluate our algorithm on various MTL benchmarks and demonstrate its superior performance over existing methods in the presence of label noise.
00footnotetext: Correspondence to: Yifei He $<$yifeihe3@illinois.edu$>$, Han Zhao $<$hanzhao@illinois.edu$>$.
@@ 1 Introduction
Multi-task learning (MTL) aims to train a single model to perform multiple related tasks (Caruana, 1997). Due to the nature of learning multiple tasks simultaneously, the problem is often tackled by aggregating multiple objectives into a scalar one via a convex combination. Despite various efforts to achieve more balanced training, a crucial aspect that often remains overlooked is robustness to label noise. Label noise is ubiquitous in real-world MTL problems as the tasks are drawn from diverse sources, introducing variations in data quality (Hsieh and Tseng, 2021; Burgert et al., 2022). The presence of label noise can negatively impact the performance of all tasks trained jointly, leading to sub-optimal performance across the board. Addressing this issue is essential for the robust and reliable deployment of MTL models in real-world scenarios.
@@
Theoretically, we derive convergence guarantees for our algorithm and establish connections with multi-objective optimization, proving that the solutions of ExcessMTL are Pareto stationary. Empirically, we evaluate our method on various MTL benchmarks, showing that it outperforms existing adaptive weighting methods in the presence of label noise, even in cases of extreme noise in one or a few tasks. Our results highlight the robustness of excess risk-based weighting methods, especially when label noise is a concern.
@@ 2.1 Excess Risks
Consider predicting the label $y\in\mathcal{Y}$ from the input data $x\in\mathcal{X}$. Let the data be drawn from some distribution $P$. Given a loss function $\ell$, the risk (or expected loss) of a model $\theta$ from the model family $\Theta$ is given by
$\varepsilon(\theta)=\mathbb{E}_{(x,y)\sim P}[\ell(\theta;(x,y))].$
@@
$\varepsilon(\theta)=\underbrace{\underbrace{\varepsilon(\theta)-\varepsilon(\theta_{\Theta}^{*})}_{\text{Estimation error}}+\underbrace{\varepsilon(\theta_{\Theta}^{*})-\varepsilon(\theta^{*})}_{\text{Approximationerror}}}_{\text{Excess risk}(\mathcal{E})}+\underbrace{\varepsilon(\theta^{*})}_{\text{Bayes error}},$
where $\theta_{\Theta}^{*}=\operatornamewithlimits{argmin}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim P}[\ell(\theta;(x,y))]$ is the optimal model in the model family $\Theta$ and $\theta^{*}=\operatornamewithlimits{argmin}_{\theta}\mathbb{E}_{(x,y)\sim P}[\ell(\theta;(x,y))]$ is the optimal model in any model family. The combination of estimation error and approximation error is the excess risk, denoted as $\mathcal{E}$. When the function class $\mathcal{F}$ is expressive enough, the approximation error approaches 0. The Bayes error is irreducible due to the stochasticity in the data generating process (e.g., label noise). For instance, if the data generating process is non-deterministic, one data point can have non-zero probability of belonging to multiple classes. This is an inherent property of a dataset, where high label noise leads to high Bayes error. On the other hand, the excess risk captures the difference between the risks of the model and the Bayes optimal model, effectively removing the influence of label noise. Therefore, it can be viewed as a measure of distance to optimality.
@@ 2.2 Multi-task Learning
In MTL, $m\geq 2$ tasks are given. We use $\alpha_{i}$ to represent the weight of the $i$th task and $\Delta_{m}$ to denote the $(m-1)$-dimensional probability simplex. We study the setting of hard parameter sharing, where a subset of model parameters ($\theta_{sh}$) is shared across all tasks, while other parameters ($\theta_{i}$) are task-specific. The goal of MTL is to find the parameter $\theta_{sh}$ and $\theta_{i}$ that minimizes a convex combination of all task-specific losses ($\ell_{i}$)
$\min_{\theta_{sh},\theta_{1},\cdots\theta_{m}}\sum_{i=1}^{m}\alpha_{i}\ell_{i}(\theta_{sh},\theta_{i}),$ (1)
@@
Despite the wide usage of the weighted combination scheme, it is difficult to define optimality under the MTL setting because a model may work well on some tasks, but perform poorly on others. For more rigorous optimality analysis, MTL can be formulated as a multi-objective optimization problem (Sener and Koltun, 2018; Zhou et al., 2022), where two models can be compared by Pareto dominance.
@@ 3.1 Motivations and Objectives
where $\mathcal{E}_{i}$ is the excess risk of the $i$th task.
@@ 3.2 Algorithm
The factor $1/2$ can be dropped for simplicity. However, the computation of the Hessian matrix is generally intractable, so we use the diagonal approximation of empirical Fisher (Amari, 1998) to estimate it, which computes a diagonal matrix through the accumulation of outer products of historical gradients. This approach has shown to be useful in various practical settings (Duchi et al., 2011; Kingma and Ba, 2015). Specifically, let $g_{i}^{(\tau)}$ be the gradient of the $i$th task with respect to the model parameter $\theta_{i}^{(\tau)}$ at time step $\tau$, the approximate Hessian at time step $t$ for the $i$th task is
@@
$\min_{\theta_{sh},\theta_{1},\cdots\theta_{m}}\max_{\alpha\in\Delta_{m}}\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\theta_{sh},\theta_{i}).$ (8)
@@
where $\eta_{\alpha}$ is the step size for weight update. Note that this update can be viewed as an exponentiated gradient (Kivinen and Warmuth, 1997) step on the convex combination of excess risks. Based on the updated task weights, we compute the loss by taking a convex combination of each task-specific loss and backpropagating the gradient. Multiplicative weight update provides stability in training and we show the convergence guarantees in Section 3.4.
@@ 3.3 Conceptual Comparison
MGDA (Sener and Koltun, 2018) formulates multi-task learning as multi-objective optimization. They solve it by using the classical multiple gradient descent algorithm (Mukai, 1980; Fliege and Svaiter, 2000; Désidéri, 2012), which finds the minimum norm within the convex hull of task gradients. It also tends to favor tasks with small gradient magnitudes due to the nature of the Frank-Wofle algorithm it employs. Consider two tasks with gradients $\|g_{1}\|_{2}>\|g_{2}\|_{2}$ where the projection of $g_{1}$ on $g_{2}$ has a larger norm than $g_{2}$. In this case, MGDA will concentrate all weights on task 2. This observation holds in general beyond the two-task example. As the loss landscape of a noisy task tends to be flatter, its gradient magnitude will be smaller, thus favored by MGDA. Moreover, the Frank-Wofle algorithm requires pairwise computation among all tasks, which is prohibitive under large number of tasks.
@@ 3.4 Theoretical Analysis
To validate the theoretical soundness of Algorithm 1, we analyze its convergence property. Algorithm 1 takes inspiration from the online mirror descent algorithm (Nemirovskij and Yudin, 1983), where the update of parameter $\theta$ and weights $\alpha$ corresponds to online gradient descent and online exponentiated gradient respectively. With established results (Nemirovski et al., 2009) in the online learning literature, we show that Algorithm 1 converges at the rate $\mathcal{O}(1/\sqrt{t})$.
@@
Suppose (i) each task-specific loss $\ell_{i}$ is L-Lipschitz, (ii) $\ell_{i}$ is convex on the model parameter $\theta$, (iii) $\ell_{i}$ bounded by $B_{\ell}$ and (iv) $\|\theta\|_{2}$ is bounded by $B_{\theta}$. At time step $t$, let $\bar{\theta}^{(1:t)}\coloneqq\frac{1}{t}\sum_{\tau=1}^{t}{\theta^{(\tau)}}$, then
$\mathbb{E}\left[\sum_{i=1}^{m}\alpha_{i}^{(t)}\mathcal{E}_{i}(\bar{\theta}^{(1:t)})\right]-\min_{\theta}\max_{\alpha\in\Delta_{m}}\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\theta)\leq 2m\sqrt{\frac{10(B_{\theta}^{2}L^{2}+B_{\ell}^{2}\log m)}{t}},$
@@
With the convergence analysis, we further deduce that the solution of Algorithm 1 is Pareto optimal in the convex setting.
@@
$\mathbb{E}\left[\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\bar{\theta}^{(1:t)})\right]-\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\theta^{*})\leq$ $2m\sqrt{\frac{10(B_{\theta}^{2}L^{2}+B_{\ell}^{2}\log m)}{t}},$
@@
Suppose each task-specific loss $\ell_{i}$ is (i) L-Lipschitz (ii) G-Smooth and (iii) bounded by $B_{\ell}$. At time step $t$, using the weights $\alpha$ output by Algorithm 1, we have
@@
where $m$ is the number of tasks.
This shows that Algorithm 1 converges to a Pareto stationary point with the rate $\mathcal{O}(1/t^{1/4})$, matching the rate for single-objective SGD (Drori and Shamir, 2020). We also empirically show that the algorithm performs well under the non-convex setting in Section 4. The proof is in Appendix B.
@@ 5 Related Work
MTL Optimization can be tackled from multiple perspectives. One line of work focuses on network architecture design (Dai et al., 2016; Liu et al., 2019). Another approach is gradient manipulation, where Chen et al. (2020) drops gradient components by the extent of conflict; Yu et al. (2020) removes gradient conflict by projection; Liu et al. (2021a) minimizes average loss while tracking the improvement on the worst task. Researchers have also studied the problem from a game theory perspective (Navon et al., 2022). In this study, our primary focus lies on task weighting methods (Kendall et al., 2018; Chen et al., 2018; Sener and Koltun, 2018; Liu et al., 2021c; Lin et al., 2022).
Task weighting is a common strategy employed to prioritize tasks with inferior performance. Existing algorithms in this domain differ in their measures of task difficulty, including homoscedastic uncertainty (Kendall et al., 2018), task-specific performance measures (Guo et al., 2018), norms of gradients and training rates (Chen et al., 2018), or a combination of task-specific losses and gradients (Liu et al., 2021c). Another approach formulates the MTL problem as a multi-objective optimization problem (Sener and Koltun, 2018) and utilizes the classical multiple-gradient descent algorithm (MGDA) (Mukai, 1980; Fliege and Svaiter, 2000; Désidéri, 2012) to find Pareto stationary solutions. While MGDA does not explicitly use gradient norms for task weighting, it implicitly considers them by seeking the minimum norm point within the convex hull of task gradients.
@@ Acknowledgements
This work is partially supported by a research grant from the Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE) and AWS Cloud Credits.
@@ Appendix A Proof for Pareto Optimality
We directly apply the well-established regret bound for online mirror descent from Nemirovski et al. [2009] on the saddle point problem
@@
In our case, $f_{i}(\cdot)$ is the excess risk for the $i$th task $\mathcal{E}_{i}(\cdot)$. We take inspiration from the proof strategy of Proposition 2 in Sagawa et al. [2020] to first present the bound and then explain why our formulation satisfies all conditions for the bound.
@@
If Assumption 1-3 hold, at time step $t$, the regret for the online mirror descent algorithm on the saddle point problem

### v3

@@ Abstract
Multi-task learning (MTL) considers learning a joint model for multiple tasks by optimizing a convex combination of all task losses. To solve the optimization problem, existing methods use an adaptive weight updating scheme, where task weights are dynamically adjusted based on their respective losses to prioritize difficult tasks. However, these algorithms face a great challenge whenever label noise is present, in which case excessive weights tend to be assigned to noisy tasks that have relatively large Bayes optimal errors, thereby overshadowing other tasks and causing performance to drop across the board. To overcome this limitation, we propose Multi-Task Learning with Excess Risks (ExcessMTL), an excess risk-based task balancing method that updates the task weights by their distances to convergence instead. Intuitively, ExcessMTL assigns higher weights to worse-trained tasks that are further from convergence. To estimate the excess risks, we develop an efficient and accurate method with Taylor approximation. Theoretically, we show that our proposed algorithm achieves convergence guarantees and Pareto stationarity. Empirically, we evaluate our algorithm on various MTL benchmarks and demonstrate its superior performance over existing methods in the presence of label noise. Our code is available at https://github.com/yifei-he/ExcessMTL.
Machine Learning, ICML
@@ 1 Introduction
Multi-task learning (MTL) aims to train a single model to perform multiple related tasks (Caruana, 1997). Due to the nature of learning multiple tasks simultaneously, the problem is often tackled by aggregating multiple objectives into a scalar one via a convex combination. Despite various efforts to achieve more balanced training, a crucial aspect that often remains overlooked is robustness to label noise. Label noise is ubiquitous in real-world MTL problems as the tasks are drawn from diverse sources, introducing variations in data quality (Hsieh & Tseng, 2021; Burgert et al., 2022). The presence of label noise can negatively impact the performance of all tasks trained jointly, leading to sub-optimal performance across the board. Addressing this issue is essential for the robust and reliable deployment of MTL models in real-world scenarios.
@@
Theoretically, we derive convergence guarantees for our algorithm and establish connections with multi-objective optimization, proving that the solutions of ExcessMTL are Pareto stationary. Empirically, we evaluate our method on various MTL benchmarks, showing that it outperforms existing adaptive weighting methods in the presence of label noise, even in cases of extreme noise in one or a few tasks. Our results highlight the robustness of excess risk-based weighting methods, especially when label noise is a concern. Beyond MTL, our insights on the connection between excess risks and convergence distance can potentially inspire more robust algorithmic design for applications using loss weighting methods.
@@ 2.1 Excess Risks
Consider predicting the label $y\in\mathcal{Y}$ from the input data $x\in\mathcal{X}$. Let the data be drawn from some distribution $P$. Given the per-sample loss function $\mathcal{L}$, the risk (or expected loss) of a model $\theta$ from the model family $\Theta$ is given by
$\ell(\theta)=\mathbb{E}_{(x,y)\sim P}[\mathcal{L}(\theta;(x,y))].$
@@
$\ell(\theta)=\underbrace{\underbrace{\ell(\theta)-\ell(\theta_{\Theta}^{*})}_{\text{Estimation error}}+\underbrace{\ell(\theta_{\Theta}^{*})-\ell(\theta^{*})}_{\text{Approximation error}}}_{\text{Excess risk}(\mathcal{E})}+\underbrace{\ell(\theta^{*})}_{\text{Bayes error}},$
where $\theta_{\Theta}^{*}=\operatornamewithlimits{argmin}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim P}[\mathcal{L}(\theta;(x,y))]$ is the optimal model in the model family $\Theta$ and $\theta^{*}=\operatornamewithlimits{argmin}_{\theta}\mathbb{E}_{(x,y)\sim P}[\mathcal{L}(\theta;(x,y))]$ is the optimal model in any model family. The combination of estimation error and approximation error is the excess risk, denoted as $\mathcal{E}$. When the function class $\mathcal{F}$ is expressive enough, the approximation error approaches 0. The Bayes error is irreducible due to the stochasticity in the data generating process (e.g., label noise). For instance, if the data generating process is non-deterministic, one data point can have non-zero probability of belonging to multiple classes. This is an inherent property of a dataset, where high label noise leads to high Bayes error. On the other hand, the excess risk captures the difference between the risks of the model and the Bayes optimal model, effectively removing the influence of label noise. Therefore, it can be viewed as a measure of distance to optimality.
@@ 2.2 Multi-Task Learning
In MTL, $m\geq 2$ tasks are given. We use $\alpha_{i}$ to represent the weight of the $i$-th task and $\Delta_{m}$ to denote the $(m-1)$-dimensional probability simplex. We study the setting of hard parameter sharing, where a subset of model parameters ($\theta_{sh}$) is shared across all tasks, while other parameters ($\theta_{i}$) are task-specific. The goal of MTL is to find the parameter $\theta_{sh}$ and $\theta_{i}$ that minimizes a convex combination of all task-specific losses ($\ell_{i}$)
$\min_{\theta_{sh},\theta_{1},\cdots,\theta_{m}}\sum_{i=1}^{m}\alpha_{i}\ell_{i}(\theta_{sh},\theta_{i}),$ (1)
@@
Despite the wide usage of the weighted combination scheme, it is difficult to define optimality under the MTL setting because a model may work well on some tasks, but perform poorly on others. For more rigorous optimality analysis, MTL can be formulated as a multi-objective optimization problem (Sener & Koltun, 2018; Zhou et al., 2022), where two models can be compared by Pareto dominance.
@@ 3.1 Motivations and Objectives
where $\mathcal{E}_{i}$ is the excess risk of the $i$-th task. Note that the formulation has the same form as Tchebycheff scalarization function, whose solution is able to fully explore the Pareto front of multiple objectives (Bowman Jr, 1976; Steuer & Choo, 1983), a property lacking in linear scalarization (Hu et al., 2023).
@@ 3.2 Algorithm
The factor $1/2$ can be dropped for simplicity. However, the computation of the Hessian matrix is generally intractable, so we use the diagonal approximation of empirical Fisher (Amari, 1998) to estimate it, which computes a diagonal matrix through the accumulation of outer products of historical gradients. This approach has shown to be useful in various practical settings (Duchi et al., 2011; Kingma & Ba, 2015). Specifically, let $g_{i}^{(\tau)}$ be the gradient of the $i$-th task with respect to the model parameter $\theta_{i}^{(\tau)}$ at training step $\tau$, the approximate Hessian at training step $t$ for the $i$-th task is
@@
$\min_{\theta_{sh},\theta_{1},\cdots,\theta_{m}}\max_{\alpha\in\Delta_{m}}\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\theta_{sh},\theta_{i}).$ (8)
@@
where $\eta_{\alpha}$ is the step size for weight update. Note that this update can be viewed as an exponentiated gradient (Kivinen & Warmuth, 1997) step on the convex combination of excess risks. Based on the updated task weights, we compute the loss by taking a convex combination of each task-specific loss and backpropagating the gradient. Multiplicative weight update provides stability in training and we show the convergence guarantees in Section 3.4.
@@ 3.3 Conceptual Comparison
MGDA (Sener & Koltun, 2018) formulates multi-task learning as multi-objective optimization. They solve it by using the classical multiple gradient descent algorithm (Mukai, 1980; Fliege & Svaiter, 2000; Désidéri, 2012), which finds the minimum norm within the convex hull of task gradients. It also tends to favor tasks with small gradient magnitudes due to the nature of the Frank-Wofle algorithm it employs. Consider two tasks with gradients $\|g_{1}\|_{2}>\|g_{2}\|_{2}$ where the projection of $g_{1}$ on $g_{2}$ has a larger norm than $g_{2}$. In this case, MGDA will concentrate all weights on task 2. This observation holds in general beyond the two-task example. As the loss landscape of a noisy task tends to be flatter, its gradient magnitude will be smaller, thus favored by MGDA. Moreover, the Frank-Wofle algorithm requires pairwise computation among all tasks, which is prohibitive under large number of tasks.
@@ 3.4 Theoretical Analysis
In this section, we analyze ExcessMTL theoretically, and verify its soundness with three key properties: ExcessMTL i) converges at the same rate as MGDA and GroupDRO (Theorem 3.1), ii) converges to Pareto optimal solutions in convex cases (Corollary 3.2), iii) converges to Pareto stationary solutions in non-convex cases (Theorem 3.3).
Algorithm 1 takes inspiration from the online mirror descent algorithm (Nemirovskij & Yudin, 1983), where the update of parameter $\theta$ and weights $\alpha$ corresponds to online gradient descent and online exponentiated gradient respectively. With established results (Nemirovski et al., 2009) in the online learning literature, we show that Algorithm 1 converges at the rate $\mathcal{O}(1/\sqrt{t})$, same as MGDA and GroupDRO.
@@
Suppose (i) each task-specific loss $\ell_{i}$ is L-Lipschitz, (ii) $\ell_{i}$ is convex on the model parameter $\theta$, (iii) $\ell_{i}$ bounded by $B_{\ell}$ and (iv) $\|\theta\|_{2}$ is bounded by $B_{\theta}$. At training step $t$, let $\bar{\theta}^{(1:t)}\coloneqq\frac{1}{t}\sum_{\tau=1}^{t}{\theta^{(\tau)}}$, then
$\mathbb{E}\left[\sum_{i=1}^{m}\alpha_{i}^{(t)}\mathcal{E}_{i}(\bar{\theta}^{(1:t)})\right]-\min_{\theta}\max_{\alpha\in\Delta_{m}}\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\theta)$
$\leq$ $2m\sqrt{\frac{10(B_{\theta}^{2}L^{2}+B_{\ell}^{2}\log m)}{t}},$ (10)
@@
With the convergence analysis, we further deduce that the solution of ExcessMTL is Pareto optimal in the convex setting.
@@
$\mathbb{E}\left[\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\bar{\theta}^{(1:t)})\right]-\sum_{i=1}^{m}\alpha_{i}\mathcal{E}_{i}(\theta^{*})$
$\leq$ $2m\sqrt{\frac{10(B_{\theta}^{2}L^{2}+B_{\ell}^{2}\log m)}{t}},$ (11)
@@
Suppose each task-specific loss $\ell_{i}$ is (i) L-Lipschitz (ii) G-Smooth and (iii) bounded by $B_{\ell}$. At training step $t$, using the weights $\alpha$ output by Algorithm 1, we have
@@
where $m$ is the number of tasks, and $\sigma$ corresponds to the assumption of a bounded approximation error of the subgradient and the gradient of the excess risk detailed in A.3 of Appendix A.
This shows that ExcessMTL converges to a Pareto stationary point with the rate $\mathcal{O}(1/t^{1/4})$, matching the rate for single-objective SGD (Drori & Shamir, 2020). We also empirically show that the algorithm performs well under the non-convex setting in Section 4. The proof is in Appendix B.
The results demonstrate the theoretical soundness of our method, as it provably converges to optimal solutions in multi-objective optimization. MGDA and GroupDRO also provably converge to Pareto optimal solutions in the convex setting, while GradNorm and IMTL lack this property, meaning that their solutions can be Pareto dominated.
@@ 5 Related Work
MTL Optimization can be tackled from multiple perspectives. One line of work focuses on network architecture design (Dai et al., 2016; Liu et al., 2019). Another approach is gradient manipulation, where Chen et al. (2020) drops gradient components by the extent of conflict; Yu et al. (2020) removes gradient conflict by projection; Liu et al. (2021a) minimizes average loss while tracking the improvement on the worst task. Researchers have also studied the problem from a game theory perspective (Navon et al., 2022). In this study, our primary focus lies on task weighting methods (Kendall et al., 2018; Chen et al., 2018; Sener & Koltun, 2018; Liu et al., 2021c; Lin et al., 2022).
Task weighting is a common strategy employed to prioritize tasks with inferior performance. Existing algorithms in this domain differ in their measures of task difficulty, including homoscedastic uncertainty (Kendall et al., 2018), task-specific performance measures (Guo et al., 2018), norms of gradients and training rates (Chen et al., 2018), or a combination of task-specific losses and gradients (Liu et al., 2021c). Another approach formulates the MTL problem as a multi-objective optimization problem (Sener & Koltun, 2018) and utilizes the classical multiple-gradient descent algorithm (MGDA) (Mukai, 1980; Fliege & Svaiter, 2000; Désidéri, 2012) to find Pareto stationary solutions. While MGDA does not explicitly use gradient norms for task weighting, it implicitly considers them by seeking the minimum norm point within the convex hull of task gradients.
@@ Acknowledgements
HZ is partially supported by a research grant from the Amazon-Illinois Center on AI for Interactive Conversational Experiences (AICE) and a Google Research Scholar Award.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proof for Pareto Optimality
We directly apply the well-established regret bound for online mirror descent from Nemirovski et al. (2009) on the saddle point problem
@@
In our case, $f_{i}(\cdot)$ is the excess risk for the $i$-th task $\mathcal{E}_{i}(\cdot)$. We take inspiration from the proof strategy of Proposition 2 in (Sagawa et al., 2020) to first present the bound and then explain why our formulation satisfies all conditions for the bound.
@@
If Assumption 1-3 hold, at training step $t$, the regret for the online mirror descent algorithm on the saddle point problem